(2) Determine the truth value of each of these statements if the domain of each variable consists of all real numbers.
(a)∀x∃y(x^2=y)
(b)∀x∃y(x=y^2)
(c)∃x∀y(xy= 0)
(d)∃x∃y(x+y=/=y+x)
(e)∀x(x=/= 0 => ∃y(xy= 1))
(f)∃x∀y(y=/= 0 => xy= 1)
(g)∀x∃y(x+y= 1)
(h)∃x∃y(x+ 2y= 2^2x+ 4y= 5)
(i)∀x∃y(x+y= 2^2xy= 1)
(j)∀x∀y∃z(z= (x+y)=2)
1
Expert's answer
2020-09-08T17:09:07-0400

Here given,

The domain of each variable consists of all real numbers.

"(a)\u2200x\u2203y(x^2=y)"


For every real value of x function x2 =y exist means there exist value y which also a

real number.So truth value = TRUE.



"(b)\u2200x\u2203y(x=y^2)"


function is x=y2 which is not true for every x . if x=(-1) then not exist y which square value is (-1) . So, truth value = FALSE.


"(c)\u2203x\u2200y(xy= 0)"


function is xy = 0. here for every value of y there exist x that is x=0. so, xy=0.

So truth value = TRUE.


"(d)∃x∃y(x+y {=}\mathllap{/\,} y+x)"

function is "x+y {=}\mathllap{/\,} y+x" is false because addition of real number is commutative.

So, truth value = FALSE.


"(e)∀x(x{=}\mathllap{/\,}0\implies ∃y(xy= 1))"

for any nonzero x, the assignment y = 1/x has the property that is xy = x.(1/x) = 1.

So truth value = TRUE.


"(f)∃x∀y(y{=}\mathllap{/\,} 0 \implies xy= 1)"


This is false. If it were true, then there would be an x for which the proposition is true for

y = 2 and 3 simultaneously, i.e., 2x = 1 and 3x = 1. Taking the difference of these

two equations, we find x = 3x − 2x = 1 − 1 = 0,


but x · y = 0 · y = 0 for any y, making our premise absurd.


"(g)\u2200x\u2203y(x+y= 1)"


function is "x+y =1" is true when we take y = 1-x . so there exist value of y which make

x+y =1.


So, truth value TRUE.



"(h)∃x∃y(x+ 2y= 2 \land 2x+ 4y= 5)"


here 2x+4y = 5 divide this eqution by 2 then we get the eqution x+2y = 5/2;

compare with line eqution y = mx+c

The line it determines is of the same slope as the first equation, but they have distinct

y-intercepts, which means that they are parallel and so, never intersect. So, they

cannot have a common solution (x, y). 

So, truth value TRUE.


"(i)∀x∃y(x+y= 2 \land xy= 1)"


if x= 1 and y=1 then this is solution that satifies given condition of eqution .

So, truth value TRUE.


"(j)\u2200x\u2200y\u2203z(z= (x+y)=2)"


here x+y = 2 means that we can take y = 2-x; so real of x and y are exist that gives z=2;


So, truth value TRUE.